A 3.1g sample of unknown organic gas molecules composed of carbon, hydrogen and oxygen undergoes complete combustion to produce a 4.4g of carbon dioxide and 2.7g of water. Calculate the empirical and molecular formula of the compound if the molar mass is 62g/mol. Show working.


1
Expert's answer
2021-04-25T03:48:42-0400

mass of CO2 = 4.4 g , molar mass of CO2 = 44g , mass of H2O =2.7 g, molar mass of H2O = 18 g

mass of unknown organic gas = 3.1 g


moles of CO2 = "\frac{mass}{molar mass}= \frac{4.4}{44}= 0.1"


moles of H2O = "\frac{mass}{molar mass}= \frac{2.7}{18}=0.15"


mass of oxygen = 4.4 + 2.7 -3.1 =4.0 g


mole of O2 = "\frac{mass}{molar mass}= \frac{4.0}{32}= 0.125"


so, ratio of O2: CO2: H2O= 0.125:0.1:0.15

O2: CO2: H2O= 5:4:6


so we can write the balanced equation as


"nC_xH_yO_z + 5 O_2 \longrightarrow 4CO_2 + 6H_2O"

from above balanced equation we can conclude that ratio in which C,H and O are present

C:H:O = 4:12:4

C:H:O = 1:3:1

so, empirical formula is CH3O , and empirical mass = 31

since we are gettting empirical mass as 31 g/mol whereas molar mass = 62 g/mol

so, n= "\frac{62}{31}=2"

Molecular formula of compound = C2H6O2